# Solving Nonlinear PDEs with Sparse Radial Basis Function Networks

### v1

@@
Abstract.
We propose a novel framework for solving nonlinear PDEs using sparse radial basis function (RBF) networks. Sparsity-promoting regularization is employed to prevent over-parameterization and reduce redundant features. This work is motivated by longstanding challenges in traditional RBF collocation methods, along with the limitations of physics-informed neural networks (PINNs) and Gaussian process (GP) approaches, aiming to blend their respective strengths in a unified framework. The theoretical foundation of our approach lies in the function space of Reproducing Kernel Banach Spaces (RKBS) induced by one-hidden-layer neural networks of possibly infinite width. We prove a representer theorem showing that the solution to the sparse optimization problem in the RKBS admits a finite solution and establishes error bounds that offer a foundation for generalizing classical numerical analysis. The algorithmic framework is based on a three-phase algorithm to maintain computational efficiency through adaptive feature selection, second-order optimization, and pruning of inactive neurons. Numerical experiments demonstrate the effectiveness of our method and highlight cases where it offers notable advantages over GP approaches. This work opens new directions for adaptive PDE solvers grounded in rigorous analysis with efficient, learning-inspired implementation.
Key words and phrases:
Sparse RBF networks; Nonlinear PDEs; Reproducing Kernel Banach Spaces; Representer theorem; Adaptive feature selection; Convergence analysis; Adaptive collocation solver
1. Introduction
@@
(1.2) $\mathcal{E}[u](x):=E(x,u(x),\nabla u(x),\nabla^{2}u(x)),\quad\mathcal{B}[u](x):=B(x,u(x),\nabla u(x)).$
@@
Classical numerical methods for solving PDEs, such as finite difference methods and finite element methods, have been extensively developed and broadly applied since the mid-20th century. Meshless methods represent a more recent development, with origins tracing back at least to the 1970s through the introduction of smoothed particle hydrodynamics for astrophysical simulations [27, 37]. These methods gained increasing attention in the 1990s, leading to the emergence of various approaches, including the element-free Galerkin method, reproducing kernel particle methods, radial basis function (RBF) methods, and others [6, 11, 23, 24, 33, 36, 61]. A notable connection exists between meshless methods and machine learning through the use of kernel techniques, which have achieved considerable success in various machine learning applications [53]. The link between kernel-based approaches in machine learning and numerical solutions of PDEs was insightfully highlighted in the survey by Schaback and Wendland (2006) [52]. Kernel-based approaches are particularly effective for linear problems, such as regression or numerical solutions of linear PDEs, where the underlying solvers depend on techniques from numerical linear algebra. The recent popularity of neural networks has shifted the focus toward inherently nonlinear techniques, even for linear problems [21, 28, 47]. As a result, optimization has begun to replace traditional linear algebra as the central computational framework. However, neural network based approaches, such as physics-informed neural networks (PINNs) [47], present significant challenges in training, due to their sensitivity to hyperparameter tuning and the substantial computational cost associated with large, often over-parameterized architectures that may include many redundant features. A notable recent development is the Gaussian process (GP) approaches [5, 15, 16], which extend kernel-based methods to solving nonlinear PDEs. The central idea of GP typically involves solving a linear system as an inner step combined with an outer optimization step. However, since they are reduced to classical kernel-based or RBF methods in linear cases, many of the limitations associated with traditional RBF approaches still apply.
@@
This work is originally motivated by the aim of addressing some of the persistent limitations of traditional RBF collocation methods. While RBF collocation methods are well known for their advantages, such as being inherently meshfree and offering fast or even spectral convergence in certain cases, they also face significant challenges. A central difficulty lies in the choice of the kernel scaling parameter, which plays a critical role in the success of the method but lacks a universal selection criterion. Choosing this parameter typically involves a trade-off between approximation accuracy and the condition number of the resulting system, and there is no clear guideline for achieving an optimal balance. For large linear systems, RBF collocation often leads to extremely ill-conditioned systems. In their 2006 survey, Schaback and Wendland [52] envisioned that one important direction for developments in kernel-based approximation should focus on adaptive algorithms that solve the problems approximately, thereby alleviating the issue of ill-conditioning in large systems. The proposed approach, which is broadly applicable, is an actualization of their concept of adaptivity by selecting only a small number of features through sparsity-promoting regularization. Importantly, the kernel scale parameter is also determined adaptively, eliminating the need for tuning and resolving the long-standing issue of scale selection. This represents a new development, even in the context of kernel-based linear regression or function approximation. Our work is also motivated by the goal of addressing some of the issues of PINNs and GP methods. In this sense, the proposed approach can be viewed as: (1) a PINN framework trained using a shallow and sparse neural network for more efficient optimization, and (2) an extension of GP methods that enables adaptive selection of kernel parameters. We note that the proposed approach can be viewed as a relaxation of kernel collocation methods, as the empirical loss is constructed from pointwise evaluations of PDEs. This leads to an interpretation of the method as an “adaptive collocation solver”. The formulation is intentionally simple, allowing us to clearly demonstrate the potential of the idea, both in terms of theoretical foundation and practical performance. Lastly, the adaptivity in our approach lies in the choice of trial functions. Other extensions, such as alternative loss function formulations or adaptivity in the selection of test points, may serve as promising directions for future research.
@@
1.1. Summary of approach and contributions
@@
(1.3) $u_{c,\omega}(x)=\sum_{n=1}^{N}c_{n}\varphi(x;\omega_{n}),$
@@
(1.4) $L(u)$ $=\frac{1}{2}\lVert\mathcal{E}[u]\rVert^{2}_{L^{2}(\nu_{D})}+\frac{\lambda}{2}\lVert\mathcal{B}[u]\rVert^{2}_{L^{2}(\nu_{\partial D})}$
@@
(1.5) $\hat{L}(u)=\frac{1}{2}\sum_{k=1}^{K_{1}}w_{1,k}(\mathcal{E}[u](x_{1,k}))^{2}+\frac{\lambda}{2}\sum_{k=1}^{K_{2}}w_{2,k}(\mathcal{B}[u](x_{2,k}))^{2},$
@@
($P_{N}^{\text{emp}}$) $\min_{N,c,\omega}\hat{L}(u_{c,\omega})+\alpha\|c\|_{1}$
Unlike most existing approaches, we do not fix the network width $N$ a prior; instead, it is treated as part of the optimization problem. However, under this formulation, the existence of a minimizer for the empirical problem ($P_{N}^{\text{emp}}$) is unclear. Conceptually, without a bound on the network width $N$, we might get a network solution of infinite width. Hence, we generalize the discrete network representation (1.3) to a continuous formulation by introducing the integral neural network [2, 7, 44, 49], where $u$ is defined by
@@
Here $M(\Omega)$ is the space of signed Radon measure endowed with total variation norm $\|\cdot\|_{M(\Omega)}$. Therefore, we consider the continuous optimization problems
($P_{\mu}$) $\min_{\mu\in M(\Omega)}L(u_{\mu})+\alpha\|\mu\|_{M(\Omega)}$
($P_{\mu}^{\text{emp}}$) $\min_{\mu\in M(\Omega)}\hat{L}(u_{\mu})+\alpha\|\mu\|_{M(\Omega)}$
@@
We will establish the existence of minimizers for both ($P_{\mu}$) and ($P_{\mu}^{\text{emp}}$). More significantly, we will show that ($P_{\mu}^{\text{emp}}$) possesses a finite solution, reducing the problem to the discrete formulation ($P_{N}^{\text{emp}}$). Nevertheless, solving ($P_{N}^{\text{emp}}$) remains nontrivial since $N$ cannot be optimized through gradient-based methods. We then employ an adaptive search of neurons, where neurons are dynamically added and removed to maintain a compact network structure while minimizing the regularized loss. It follows an iterative refinement process found in sparse approximation and boosting-style methods [25, 40]. The results can accurately approximate the PDE solution with high sparsity.
@@
This is a reproducing kernel Banach space (RKBS) by definition in [4], and $\varphi:D\to C_{0}(\Omega)$ is the associated feature map. Equivalently, ($P_{\mu}$) and ($P_{\mu}^{\text{emp}}$) can be reformulated as
(1.6) $\min_{u\in\mathcal{V}(D)}L(u)+\alpha\|u\|_{\mathcal{V}}\quad\text{and}\quad\min_{u\in\mathcal{V}(D)}\hat{L}(u)+\alpha\|u\|_{\mathcal{V}}.$
These formulations identify the natural function space for shallow neural networks, where regularization is imposed through the Banach norm. For related studies on RKBS-based learning, we refer the readers to [31, 41, 42] and the references therein.
@@
    A framework integrating ideas from PINNs and GP methods: Our approach can be viewed as a PINN-type training scheme using a shallow, sparse neural network for improved optimization and an extension of the GP mechanism with adaptive kernel parameters selection.
@@
To illustrate the effectiveness of our approach, we present a simple 1D example comparing our results and the results obtained with the GP method. Figure 1(a) demonstrates that our method accurately recovers the solution using a small number of adaptively chosen kernels. Here, the stars indicate the RBF center and shape parameter, with larger stars corresponding to larger RBF bandwidths. In contrast, Figures 1(b) and 1(c) show that the GP method is highly sensitive to the kernel shape parameter. For this example, accurate recovery using GP is only possible within a very narrow range of $\sigma$ and deviations lead to significant errors.
Figure 1. Numerical experiments on a 1D semilinear Poisson equation with Sparse RBFNet (our) and Gaussian Process. The exact solution $u$ is set to be $u(x)=\tanh(10(x+0.2))-\tanh(10(x+1.0))+0.5\left(\tanh(30(x-0.4))-\tanh(30(x-0.3))\right)$. $K=100$ grid collocation points are used for both methods; $L^{2}$ and $L^{\infty}$ error shown above are estimated at test grid of size $200$. (A) Sparse RBFNet solution with $\alpha=10^{-4}$, $\lambda=100$; (B) Error curves for GP solutions with different kernel shape parameters; (C) GP solutions with different kernel shape parameters.
1.2. Related studies
@@
1.2.1. The Hilbert space setting
@@
We also observe that $\mathcal{H}$ is a reproducing kernel Hilbert space (RKHS) associated with the kernel [2]
(1.7) $k(x,x^{\prime})=\int_{\Omega}\varphi(x;\omega)\varphi(x^{\prime};\omega)\mathop{}\!\mathrm{d}\rho(\omega).$
@@
(1.8) $\min_{c\in L^{2}(\Omega,\mathop{}\!\mathrm{d}\rho)}\hat{L}(u_{c\mathop{}\!\mathrm{d}\rho})+\frac{\alpha}{2}\lVert c\rVert^{2}_{L^{2}(\Omega,\mathop{}\!\mathrm{d}\rho)}$
@@
(1.9) $\min_{u\in\mathcal{H}}\hat{L}(u)+\frac{\alpha}{2}\lVert u\rVert^{2}_{\mathcal{H}}.$
@@
which naturally implies that $\mathcal{H}\hookrightarrow\mathcal{V}$.
1.2.2. Random feature models
@@
that approximates $k$ in Equation 1.7. For a linear PDE problem, this leads to solving a linear system, which avoids the neural network training cost. However, the random feature model does not adapt to the solution landscape and may require a large number of features. This can lead to prohibitive computational costs, especially in high-dimensional settings. For related discussions, see [46, 67, 3, 45].
1.2.3. The Gaussian process method
@@
There has been a recent surge in using the Gaussian process (GP) method for solving PDEs [5, 15, 16]. The GP method in [15] employs a Gaussian kernel with a fixed shape parameter and randomly chosen centers (which are also the collocation points). The GP method is essentially equivalent to solving (1.9) with the covariance kernel given by (1.7). Since this kernel is only indirectly related to the feature function, and also depends on the density $\mathop{}\!\mathrm{d}\rho$, we can not directly compare the kernel used for the construction of the feature function and the covariance kernel resulting from this method. However, if we consider a density $\mathop{}\!\mathrm{d}\rho$ that is concentrated on a single RBF shape parameter, but otherwise uniform over the kernel centers, we have a direct relation: Consider the case where $\omega=y\in D_{1}\subset\mathbb{R}^{d}$, $\sigma>0$ is fixed, and $\rho$ is a uniform probability measure on $D_{1}$. Then Equation 1.7 becomes
@@
$\begin{split}k(x,x^{\prime})&=\frac{\sigma^{2(s-d)}}{|D_{1}|}\int_{D_{1}}\hat{\varphi}((x-y)/\sigma)\hat{\varphi}((x^{\prime}-y)/\sigma)\mathop{}\!\mathrm{d}y\\ &\approx\frac{\sigma^{2(s-d)}}{|D_{1}|}\int_{\mathbb{R}^{d}}\hat{\varphi}((x-y)/\sigma)\hat{\varphi}((x^{\prime}-y)/\sigma)\mathop{}\!\mathrm{d}y\\ &=\frac{\sigma^{2s}}{(2\sigma)^{d}}\hat{\varphi}((x-x^{\prime})/(2\sigma))\end{split}$
@@
where $\hat{\varphi}$ denotes the standard Gaussian function in $\mathbb{R}^{d}$. This yields a Gaussian kernel with a fixed scale parameter.
1.2.4. PINNs and other machine learning method for PDEs

### v2

@@
Abstract
We propose a novel framework for solving nonlinear PDEs using sparse radial basis function (RBF) networks. Sparsity-promoting regularization is employed to prevent over-parameterization and reduce redundant features. This work is motivated by longstanding challenges in traditional RBF collocation methods, along with the limitations of physics-informed neural networks (PINNs) and Gaussian process (GP) approaches, aiming to blend their respective strengths in a unified framework. The theoretical foundation of our approach lies in the function space of Reproducing Kernel Banach Spaces (RKBS) induced by one-hidden-layer neural networks of possibly infinite width. We prove a representer theorem showing that the sparse optimization problem in the RKBS admits a finite solution and establishes error bounds that offer a foundation for generalizing classical numerical analysis. The algorithmic framework is based on a three-phase algorithm to maintain computational efficiency through adaptive feature selection, second-order optimization, and pruning of inactive neurons. Numerical experiments demonstrate the effectiveness of our method and highlight cases where it offers notable advantages over GP approaches. This work opens new directions for adaptive PDE solvers grounded in rigorous analysis with efficient, learning-inspired implementation.
Keywords: Sparse RBF networks; Nonlinear PDEs; Reproducing Kernel Banach Spaces; Representer theorem; Adaptive feature selection; Convergence analysis; Adaptive collocation solver
1 Introduction
@@
$\mathcal{E}[u](x):=E(x,u(x),\nabla u(x),\nabla^{2}u(x)),\quad\mathcal{B}[u](x):=B(x,u(x),\nabla u(x)).$ (1.2)
@@
Classical numerical methods for solving PDEs, such as finite difference methods and finite element methods, have been extensively developed and broadly applied since the mid-20th century. Meshless methods represent a more recent development, with origins tracing back at least to the 1970s through the introduction of smoothed particle hydrodynamics for astrophysical simulations (Gingold and Monaghan, 1977; Lucy, 1977). These methods gained increasing attention in the 1990s, leading to the emergence of various approaches, including the element-free Galerkin method, reproducing kernel particle methods, radial basis function (RBF) methods, and others (Belytschko et al., 1994; Buhmann, 2000; Fasshauer, 1996; Franke and Schaback, 1998; Li and Liu, 2007; Liu et al., 1996; Wendland, 2004). A notable connection exists between meshless methods and machine learning through the use of kernel techniques, which have achieved considerable success in various machine learning applications (Scholkopf and Smola, 2018). The link between kernel-based approaches in machine learning and numerical solutions of PDEs was insightfully highlighted in the survey by Schaback and Wendland (2006). Kernel-based approaches are particularly effective for linear problems, such as regression or numerical solutions of linear PDEs, where the underlying solvers depend on techniques from numerical linear algebra. The recent popularity of neural networks has shifted the focus toward inherently nonlinear techniques, even for linear problems (E and Yu, 2018; Goodfellow et al., 2016; Raissi et al., 2019). As a result, optimization has begun to replace traditional linear algebra as the central computational framework. However, neural network-based approaches, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), present significant challenges in training, due to their sensitivity to hyperparameter tuning and the substantial computational cost associated with large, often over-parameterized architectures that may include many redundant features. A notable recent development is the Gaussian process (GP) approaches (Batlle et al., 2025; Chen et al., 2021, 2025), which extend kernel-based methods to solving nonlinear PDEs. The central idea of GP typically involves solving a linear system as an inner step combined with an outer optimization step. However, since they are reduced to classical kernel-based or RBF methods in linear cases, many of the limitations associated with traditional RBF approaches still apply.
@@
This work is originally motivated by the aim of addressing some of the persistent limitations of traditional RBF collocation methods. While RBF collocation methods are well known for their advantages, such as being inherently meshfree and offering fast or even spectral convergence in certain cases, they also face significant challenges. A central difficulty lies in the choice of the kernel scaling parameter, which plays a critical role in the success of the method but lacks a universal selection criterion. Choosing this parameter typically involves a trade-off between approximation accuracy and the condition number of the resulting system, and there is no clear guideline for achieving an optimal balance. For large linear systems, RBF collocation often leads to extremely ill-conditioned systems. In their survey, Schaback and Wendland (2006) envisioned that one important direction for developments in kernel-based approximation should focus on adaptive algorithms that solve the problems approximately, thereby alleviating the issue of ill-conditioning in large systems. The proposed approach, which is broadly applicable, is an actualization of their concept of adaptivity by selecting only a small number of features through sparsity-promoting regularization. Importantly, the kernel scale parameter is also determined adaptively, eliminating the need for tuning and resolving the long-standing issue of scale selection. This represents a new development, even in the context of kernel-based linear regression or function approximation. Our work is also motivated by the goal of addressing some of the issues of PINNs and GP methods. In this sense, the proposed approach can be viewed as: (1) a PINN framework trained using a shallow and sparse neural network for more efficient optimization, and (2) an extension of GP methods that enables adaptive selection of kernel parameters. We note that the proposed approach can be viewed as a relaxation of kernel collocation methods, as the empirical loss is constructed from pointwise evaluations of PDEs. This leads to an interpretation of the method as an “adaptive collocation solver”. The formulation is intentionally simple, allowing us to clearly demonstrate the potential of the idea, both in terms of theoretical foundation and practical performance. Lastly, the adaptivity in our approach lies in the choice of trial functions. Other extensions, such as alternative loss function formulations or adaptivity in the selection of test points, may serve as promising directions for future research.
@@
1.1 Summary of approach and contributions
@@
$u_{c,\omega}(x)=\sum_{n=1}^{N}c_{n}\varphi(x;\omega_{n}),$ (1.3)
@@
$L(u)$ $=\frac{1}{2}\lVert\mathcal{E}[u]\rVert^{2}_{L^{2}(\nu_{D})}+\frac{\lambda}{2}\lVert\mathcal{B}[u]\rVert^{2}_{L^{2}(\nu_{\partial D})}$ (1.4)
@@
$\hat{L}(u)=\frac{1}{2}\sum_{k=1}^{K_{1}}w_{1,k}(\mathcal{E}[u](x_{1,k}))^{2}+\frac{\lambda}{2}\sum_{k=1}^{K_{2}}w_{2,k}(\mathcal{B}[u](x_{2,k}))^{2},$ (1.5)
@@
$\min_{N,c,\omega}\hat{L}(u_{c,\omega})+\alpha\|c\|_{1}$ ($P_{N}^{\text{emp}}$)
Unlike most existing approaches, we do not fix the network width $N$ a prior; instead, it is treated as part of the optimization problem. However, under this formulation, the existence of a minimizer for the empirical problem ($P_{N}^{\text{emp}}$) is unclear. Conceptually, without a bound on the network width $N$, we might get a network solution of infinite width. Hence, we generalize the discrete network representation (1.3) to a continuous formulation by introducing the integral neural network (Bach, 2017; Bengio et al., 2005; Pieper and Petrosyan, 2022; Rosset et al., 2007), where $u$ is defined by
@@
Here $M(\Omega)$ is the space of signed Radon measures endowed with the total variation norm $\|\cdot\|_{M(\Omega)}$. Therefore, we consider the continuous optimization problems
$\min_{\mu\in M(\Omega)}L(u_{\mu})+\alpha\|\mu\|_{M(\Omega)}$ ($P_{\mu}$)
$\min_{\mu\in M(\Omega)}\hat{L}(u_{\mu})+\alpha\|\mu\|_{M(\Omega)}$ ($P_{\mu}^{\text{emp}}$)
@@
We will establish the existence of minimizers for both ($P_{\mu}$) and ($P_{\mu}^{\text{emp}}$). More significantly, we will show that ($P_{\mu}^{\text{emp}}$) possesses a finite solution, reducing the problem to the discrete formulation ($P_{N}^{\text{emp}}$). Nevertheless, solving ($P_{N}^{\text{emp}}$) remains nontrivial since $N$ cannot be optimized through gradient-based methods. We then employ an adaptive search of neurons, where neurons are dynamically added and removed to maintain a compact network structure while minimizing the regularized loss. It follows an iterative refinement process found in sparse approximation and boosting-style methods (Friedman, 2001; Needell and Tropp, 2009). The results can accurately approximate the PDE solution with high sparsity.
@@
This is a reproducing kernel Banach space (RKBS) by definition in Bartolucci et al. (2023), and $\varphi:D\to C_{0}(\Omega)$ is the associated feature map. Equivalently, ($P_{\mu}$) and ($P_{\mu}^{\text{emp}}$) can be reformulated as
$\min_{u\in\mathcal{V}(D)}L(u)+\alpha\|u\|_{\mathcal{V}}\quad\text{and}\quad\min_{u\in\mathcal{V}(D)}\hat{L}(u)+\alpha\|u\|_{\mathcal{V}}.$ (1.6)
These formulations identify the natural function space for shallow neural networks, where regularization is imposed through the Banach norm.
The space $\mathcal{V}$ is also referred to as an integral RKBS in the literature (Spek et al., 2025), and is closely related to variation spaces and Barron spaces (E et al., 2022; Parhi and Nowak, 2021; Parhi and Unser, 2025; Siegel and Xu, 2023). For related studies on RKBS-based learning, we refer the readers to Kumar et al. (2024); Parhi and Nowak (2022); Parhi and Unser (2025) and the references therein.
@@
    A framework integrating ideas from PINNs and GP methods: Our approach can be viewed as a PINN-type training scheme using a shallow, sparse neural network for improved optimization and an extension of the GP mechanism with adaptive kernel parameter selection.
@@
To illustrate the effectiveness of our approach, we present a simple 1D example comparing our results with the results obtained with the GP method from Chen et al. (2021)  (using the code in the associated repository; Chen, 2024). Figure 1(a) demonstrates that our method accurately recovers the solution using a small number of adaptively chosen kernels. Here, the stars indicate the RBF center and shape parameter, with larger stars corresponding to larger RBF bandwidths. In contrast, Figures 1(b) and 1(c) show that the GP method is highly sensitive to the kernel shape parameter. For this example, accurate recovery using GP is only possible within a very narrow range of $\sigma$ and deviations lead to significant errors.
Figure 1: Numerical experiments on a 1D semilinear Poisson equation with Sparse RBFNet (our) and Gaussian Process. The exact solution $u$ is set to be $u(x)=\tanh(10(x+0.2))-\tanh(10(x+1.0))+0.5\left(\tanh(30(x-0.4))-\tanh(30(x-0.3))\right)$. $K=100$ grid collocation points are used for both methods; $L^{2}$ and $L^{\infty}$ error shown above are estimated at test grid of size $200$. (a) Sparse RBFNet solution with $\alpha=10^{-4}$, $\lambda=100$; (b) Error curves for GP solutions with different kernel shape parameters; (c) GP solutions with different kernel shape parameters.
1.2 Related studies
@@
1.2.1 The Hilbert space setting
@@
We also observe that $\mathcal{H}$ is a reproducing kernel Hilbert space (RKHS) associated with the kernel (Bach, 2017)
$k(x,x^{\prime})=\int_{\Omega}\varphi(x;\omega)\varphi(x^{\prime};\omega)\mathop{}\!\mathrm{d}\rho(\omega).$ (1.7)
@@
$\min_{c\in L^{2}(\Omega,\mathop{}\!\mathrm{d}\rho)}\hat{L}(u_{c\mathop{}\!\mathrm{d}\rho})+\frac{\alpha}{2}\lVert c\rVert^{2}_{L^{2}(\Omega,\mathop{}\!\mathrm{d}\rho)}$ (1.8)
@@
$\min_{u\in\mathcal{H}}\hat{L}(u)+\frac{\alpha}{2}\lVert u\rVert^{2}_{\mathcal{H}}.$ (1.9)
@@
which naturally implies that $\mathcal{H}\hookrightarrow\mathcal{V}$. This embedding is in fact strict, and Spek et al. (2025) show that $\mathcal{V}$ coincides with the union of all such RKHS as $\rho$ ranges over the set of probability measures.
1.2.2 Random feature models
@@
that approximates $k$ in Equation 1.7. For a linear PDE problem, this leads to solving a linear system, which avoids the neural network training cost. However, the random feature model does not adapt to the solution landscape and may require a large number of features. This can lead to prohibitive computational costs, especially in high-dimensional settings. For related discussions, see Rahimi and Recht (2008); Zhang et al. (2024); Bach (2023); Pieper et al. (2024).
1.2.3 The Gaussian process method
There has been a recent surge in using the Gaussian process (GP) method for solving PDEs (Batlle et al., 2025; Chen et al., 2021, 2025). The GP method proposed by Chen et al. (2021) employs a Gaussian kernel with a fixed shape parameter and randomly chosen collocation points. Solving (1.9) in the limiting case $\alpha\to 0$ yields a variant of this GP method with the covariance kernel given by (1.7), where the solution can be interpreted as a maximum likelihood estimate under a GP prior, as shown in Chen et al. (2021).
@@ 1.2.3 The Gaussian process method
However, the kernel (1.7) is only indirectly related to the feature function, and also depends on the density $\mathop{}\!\mathrm{d}\rho$, we can not directly compare the kernel used for the construction of the feature function and the covariance kernel resulting from this method. However, if we consider a density $\mathop{}\!\mathrm{d}\rho$ that is concentrated on a single RBF shape parameter, but otherwise uniform over the kernel centers, we have a direct relation: Consider the case where $\omega=y\in B_{R}\subset\mathbb{R}^{d}$ is in a ball of large radius $R\gg 0$, $\sigma>0$ is fixed, and $\rho$ is a uniform probability measure on $B_{R}$. Then Equation 1.7 becomes
@@
$\begin{split}k(x,x^{\prime})&=\frac{\sigma^{2s}}{|B_{R}|}\int_{B_{R}}\sigma^{-2d}\hat{\varphi}((x-y)/\sigma)\hat{\varphi}((x^{\prime}-y)/\sigma)\mathop{}\!\mathrm{d}y\\ &\approx\frac{\sigma^{2s}}{|B_{R}|}\int_{\mathbb{R}^{d}}\sigma^{-2d}\hat{\varphi}((x-y)/\sigma)\hat{\varphi}((x^{\prime}-y)/\sigma)\mathop{}\!\mathrm{d}y\\ &=\frac{\sigma^{2s}}{(\sqrt{2}\sigma)^{d}|B_{R}|}\hat{\varphi}\left(\frac{x-x^{\prime}}{\sqrt{2}\sigma}\right)\end{split}$
@@
where $\hat{\varphi}$ denotes the standard Gaussian function in $\mathbb{R}^{d}$. This yields a Gaussian kernel with fixed scale parameter $\sqrt{2}\sigma$.
1.2.4 PINNs and other machine learning method for PDEs